# Cross-Sectional Dynamics Under Network Structure: Theory and Macroeconomic Applications

### v6

@@ 2.3 Further Properties of the NVAR
Trivially, contemporaneous network interactions lead to contemporaneous correlation among $\{x_{i}\}_{i=1:n}$; under Eq. 6, we have $\mathbb{V}[x]=\mathbb{V}[(I-A)^{-1}v]$. In this case, $\text{Cov}(x_{i},x_{j})$ reflects the bilateral exposure between $i$ and $j$ and their mutual exposure to third units, whereby exposure is determined by network connections of all order.313131 We have $\mathbb{V}[x]=\mathbb{V}[(I+A+A^{2}+...)v]$, provided that $|\lambda_{i}|<1$ for all eigenvalues $\lambda_{i}$ of $A$. Assuming that $\mathbb{V}[v]=\Sigma=diag(\sigma^{2}_{1},...,\sigma^{2}_{n})$, we get $\text{Cov}(x_{i},x_{j})=\sigma^{2}_{i}\mathbf{1}\left\{{i=j}\right\}+\sigma^{2}_{j}\sum_{h=1}^{\infty}\left[{A^{h}}\right]_{ij}+\sigma^{2}_{i}\sum_{h=1}^{\infty}\left[{A^{h}}\right]_{ji}+\sum_{o=1}^{n}\sigma^{2}_{o}\sum_{h_{i}=1}^{\infty}\sum_{h_{j}=1}^{\infty}\left[{A^{h_{i}}}\right]_{io}\left[{A^{h_{j}}}\right]_{jo}\;.$ In the context of a production economy, this insight has been exploited for shock identification by De Graeve and Schneider (2023).
@@
Corollary 3 above establishes that an SAR for $x$ is obtained by writing $x$ as the sum of an underlying process driven by lagged network interactions that evolves at an ever higher frequency. In line with that, Proposition 3 shows that $\sqrt{q}u_{t}\overset{d}{\rightarrow}N\left({0,\mathbb{V}[(I-aA)^{-1}\tilde{u}_{\tau}]}\right)$ as $q\rightarrow\infty$ for the NVAR($p,q$) from Eq. 5 and for $a=\sum_{l=1}^{p}\alpha_{l}$ under strong stationarity of high-frequency innovations $\tilde{u}_{\tau}$. Hence, one can justify the Normality-assumption $v\sim N(0,\Sigma)$ in the SAR model by assuming temporally independent high-frequency innovations $\tilde{u}_{\tau}$ with $\mathbb{V}[\tilde{u}_{\tau}]=q^{-1}\Sigma$.
@@
The NVAR is a VAR that restricts the innovation transmission across series to occur along constant, bilateral links (in one direction). Relative to a VAR($p$) with $p>1$, the NVAR($p,1$) induces parsimony; e.g. for each $y_{it}$, the $n$ links $\{a_{ij}\}_{j=1:n}$ in $A_{i\cdot}$ summarize the information in the potentially high-dimensional set $\{y_{j,t-l}\}_{j=1:n,l=1:p}$:
@@
where $x_{it}=\sum_{j=1}^{n}a_{ij}y_{j,t-l}$ is the $i$th row of $X_{t}=[Ay_{t-1},...,Ay_{t-l}]$. If in addition the network $A$ is sparse – as is the case across a wide range of applications –, then the NVAR rationalizes the dynamic comovement among all $\{y_{it}\}_{i=1:n}$ by the dynamic innovation transmission along few bilateral links among $i=1:n$. This is akin to the assumption that longer-term dynamics are driven by a set of shorter-term dynamics, upheld by the general class of VARMA($p,q$) models.
The Dynamic Factor Model (DFM) also reduces dimensionality by summarizing a large set of predictors by a few linear combinations. Proposition 4 shows that the NVAR($p,1$) can be written as a DFM, while a DFM with restricted factor dynamics can be written as an NVAR($p,1$) in the limit as $n\rightarrow\infty$.323232 The second part of Proposition 4 restricts $f_{t}\in\mathbb{R}^{r}$ to follow an NVAR($p,1$). This is less restrictive than it may seem; for $p=1$, it requires $f_{t}\sim VAR(p)$, while under $r=1$ it requires $f_{t}\sim AR(p)$. Also, note that, as usual, the factor representation is not unique, and we can re-scale $A$ and $\phi$ to ensure $a_{ij}\in[-1,1]$. Together with Eq. 9 above, it suggests that, at the cost of restricting factor dynamics, the NVAR allows the linear combinations of predictors to vary more flexibly across units, to the point that it naturally acommodates sparse factors as (linear combinations of) locally important units in the network.333333 The DFM constructs $\{\mathbb{E}_{t-1}[y_{it}]\}_{i=1}^{n}$ as different ($\Lambda_{i\cdot}$) linear combinations of the same factors $f_{t}$, which in turn are linear combinations of observables. In contrast, the NVAR constructs $\{\mathbb{E}_{t-1}[y_{it}]\}_{i=1}^{n}$ as the same ($\alpha$) linear combination of different covariates $x_{it}$, which in turn are linear combinations of observables.
@@ 3.1 Timing of Network Effects
Section B.1 establishes consistency and asymptotic Normality of $\hat{\alpha}_{OLS}$ under large $T$, large $n$ and large $(n,T)$ asymptotics. The derivations under large $n$ assume that the observed network adjacency matrix $A_{n}$ converges to some limit $A_{*}$ so that, for example, $\frac{1}{n}\sum_{i=1}^{n}\left({A_{n,i\cdot}y_{t-l}}\right)^{\prime}u_{it}\overset{p}{\rightarrow}\mathbb{E}\left[{\left({A_{*,i\cdot}y_{t-l}}\right)^{\prime}u_{it}}\right]$.
@@
In case of an NVAR($p,q$), as defined in Eq. 4, an estimator for $\alpha$ can be obtained by data augmentation.343434 The discussion holds likewise for the NVAR($p,q$) from Eq. 5. However, point identification is not guaranteed. For example, under $q=2$ and $p=1$, the observed process follows
@@
Conditioning on $\tilde{\Sigma}$, the posterior $p(\alpha|Y_{1:T})$ can be obtained using the Gibbs sampler of Carter and Kohn (1994). Treating the unobserved data in $\tilde{Y}_{1:T_{\tau}}$ as parameters, it iteratively draws from $p(\tilde{Y}_{1:T_{\tau}}|\alpha,Y_{1:T})$ and $p(\alpha|\tilde{Y}_{1:T_{\tau}})$ to obtain a sample from $p(\alpha,\tilde{Y}_{1:T_{\tau}}|Y_{1:T})$. [^36] The particular state space model where in some periods $\tau$ no data is observed implies that in these periods the updating-step of the Kalman filter is skipped and the updated distribution of states equals their predicted distribution. See Section B.1, and see Schorfheide and Song (2015) for a discussion of the analogous case of a mixed-frequency VAR. Under a Uniform prior for $\alpha$, the resulting posterior mode of $\alpha$ converges to the ML estimator obtained using the Expectation-Maximization (EM) algorithm. To estimate $\tilde{\Sigma}$ as well, an additional iteration step is added to the Gibbs sampler. Using a uniform prior, we get an Inverse-Wishart conditional posterior for $\Sigma|\tilde{Y}_{1:T_{\tau}},\alpha$ with mode $\frac{1}{T_{\tau}}\sum_{\tau=1}^{T_{\tau}}\tilde{u}_{\tau}\tilde{u}_{\tau}^{\prime}$.
@@ 3.2 Joint Inference: Network & Effect-Timing
To render $(\alpha,A)$ jointly identified, I normalize $\alpha_{1}=1$, with appropriate redefinitions of $\alpha$ as well as $y_{t}$, $z_{t}$ and $X_{t}$. Relative to the alternative of restricting the norm of $\alpha$, this normalization facilitates analytical calculations and asymptotic analysis, but it requires $\alpha_{1}\neq 0$ in the true data generating process.373737 In case $a_{ij}\geq 0$ is restricted, it requires $\alpha_{1}>0$.
@@
Long and Plosser (1983) show that firms’ lagged conversion of inputs into outputs can generate endogenous business cycles. This contrasts with the models used in typical macroeconomic models, where business cycles are ultimately generated by persistence in exogenous shocks. Yet, Carvalho and Reischer (2021) show that the persistence obtained in the framework of Long and Plosser (1983) matches well empirical measures of persistence, suggesting that the endogenization of business cycles through lagged input-output conversion (IOC) is not only theoretically attractive, but has empirical merit as well.
This application tests the Long and Plosser (1983) hypothesis empirically. Departing from their model, I generalize the lags in IOC by assuming that firms’ production requires inputs produced in the past $p$ periods. This leads to sectoral output growth and inflation evolving at some model-frequency as an NVAR($p,1$), which translates into an NVAR($p,q$) for some $q\in\mathbb{N}$ at the monthly frequency of observation. Thereby, $A$ is the input-output matrix, $u_{t}$ contains sectoral productivity processes, while $\{\alpha_{l}\}_{l=1:p}$ show how input-sourcing is spread out over the $p$ periods. By estimating $\alpha$, $q$ and the persistence in $u_{t}$, I use the NVAR to quantify the extent to which business cycles in this framework are due to lagged IOC as opposed to persistence in exogenous productivity processes.
After theoretically motivating the analysis in Section 4.1, I discuss the application-setup, data and estimation procedure in Section 4.2. Section 4.3 discusses the preliminary results, which focus on sectoral prices and assume i.i.d. productivity shocks in order to shed light on the relation between lagged IOC and persistence in isolation. Details are in Appendix C.
@@ 4.1 Theory
The following analysis is based on Carvalho and Tahbaz-Salehi (2019), who discuss a static economy. Derivations are provided in Section C.1.
@@
where $\sum_{i=1}^{n}\gamma_{i}=1$. The first-order condition (FOC) yields $c_{i\tau}=\gamma_{i}\frac{w_{\tau}}{p_{i\tau}}$. [^40] Hence, $\gamma_{i}$ is the share of good $i$ in households’ expenditures. This result holds even if households have access to a storage technology, as market clearing under representative households in a closed economy implies that households spend their whole period $\tau$ income, $w_{\tau}$, on consumption.
@@
where $k^{y}$ is a vector of constants, and $a_{ij}=p_{j}x^{ij}/(p_{i}y_{i})$ is value of good $j$ purchases by sector $i$ as a fraction of the value of sector $i$’s output. In this environment, while the network $A$ amplifies idiosyncratic TFP shocks and therefore affects the variance of $\tilde{y}$ (see e.g. Acemoglu et al. (2012)), any autocorrelation in $\tilde{y}$ is inherited from that in $\tilde{z}$.
To analyze dynamics under lagged IOC, I assume perfect foresight.414141 As discussed in Fan et al. (2023), this assumption is standard for modeling dynamic spatial economies, which are closely related to dynamic network economies. If, as in Long and Plosser (1983), it takes one period to convert purchased inputs into output, then $x_{ij\tau}=x^{ij}_{\tau-1}$ and sectoral output approximately follows an NVAR($1,1$):
@@
where the time variation in $k^{y1}_{\tau}$ is due to time variation in the numéraire $w_{\tau}$. So long as $\beta<1$, in steady state, prices are higher and output is lower than in the economy with contemporaneous IOC. Also, the meaning of $a_{ij}=\beta^{-1}p_{j}x^{ij}/(p_{i}y_{i})$ changes slightly. These differences vanish as $\beta\rightarrow 1$. More importantly, while the former, static economy is always in steady state, the economy with one period-lagged IOC features transition dynamics; after a disturbance to $\tilde{z}_{\tau}$, $\tilde{y}_{\tau}$ only asymptotically converges to steady state.424242 Relatedly, as discussed in Section 2.2, the response of $\tilde{y}_{(\tau)}$ to a change in $\tilde{z}_{(\tau)}$ in the static economy corresponds to the long-term response of $\tilde{y}_{\tau}$ to a permanent change in $\tilde{z}_{\tau}$ in this dynamic economy (disregarding the slightly changed meaning of $a_{ij}$). It is due to these transition dynamics that this framework can generate endogenous business cycles, i.e. autocorrelation in $\tilde{y}_{\tau}$ even in absence of autocorrelation in $\tilde{z}_{\tau}$.
@@
where $x^{ij}_{\tau,\tau-h}$ denotes the use of good $j$ purchased at time $\tau-h$ in the production of good $i$ at time $\tau$.434343 This means that a good perishes after two periods (with regard to its suitability as an input in production). Therefore, the amount of good $j$ purchased at time $\tau$ can be used in production at periods $\tau+1$ and $\tau+2$: $x^{ij}_{\tau}=x^{ij}_{\tau+1,\tau}+x^{ij}_{\tau+2,\tau}$. This shortcut stands for frictions like storage capacity constraints and time-varying input delivery times.444444 As in the Long and Plosser (1983)-economy above, the presumption is that storage is done by the buyer. In the Cobb-Douglas case $r\rightarrow 0$, sectoral output approximately follows an NVAR(2,1):
@@
where once again $k^{y2}_{\tau}$ varies over time only to the extent that the numéraire $w_{\tau}$ changes in value. Under a more general elasticity of substitution $r\in[0,1)$,454545 This notably excludes complementarity ($r<0$) and perfect substitutability ($r=1$). the analogous result is obtained by log-linearizing around the steady state (see Section C.1). This specification nests the one period-lagged economy, which is obtained under $\alpha_{2}=0$. Relative to that case, $\alpha_{2}>0$ increases prices and decreases output in steady state, provided that $\beta<1$. Also, we have $a_{ij}=\left[{\beta\alpha_{1}+\beta^{2}\alpha_{2}}\right]^{-1}(p_{j}x^{ij})/(p_{i}y_{i})$. As $\beta\rightarrow 1$, the links $a_{ij}$ retain their interpretation as the output shares of different inputs $j$ in the production of good $i$. The parameters $(\alpha_{1},\alpha_{2})$ show the shares of an input $j$ purchased at different periods in the past in the overall usage of input $j$ in the production of good $i$. Their homogeneity means that the time profile of input-sourcing is assumed to be constant over time and across input-output pairs $(i,j)$. The restrictions $\alpha_{1},\alpha_{2}\geq 0$, $\alpha_{1}+\alpha_{2}=1$ and $\sum_{j=1}^{n}a_{ij}<1\;\forall\;i$ imply that $\tilde{y}_{\tau}$ is stationary as long as $\tilde{z}_{\tau}$ is.464646 Berman and Plemmons (1979, p. 37) show that for an element-wise nonnegative matrix with row sums strictly smaller than 1, the absolute value of the largest eigenvalue is strictly less than 1. Stationarity then follows by Corollary 1.
The theory above does not restrict the process of (log) sectoral TFP $\tilde{z}_{i\tau}$. In general, it may display secular growth and persistent shocks of both aggregate and idioyncratic nature. Consider a difference-stationary specification with aggregate and idiosyncratic TFP processes $e^{a}_{\tau}$ and $e_{i\tau}$ evolving as AR(1) processes:
@@
We obtain IRFs $\partial\Delta\tilde{y}_{i,\tau+h}/\partial\varepsilon_{j,\tau}$ and $\partial\Delta\tilde{y}_{i,\tau+h}/\partial\varepsilon^{a}_{\tau}$ akin to those in Eq. 7 in Section 2.3. This points to two sources of persistence in output growth: persistence in the aggregate TFP process $e^{a}_{\tau}$ and persistence in idiosyncratic TFP processes $\{e_{i\tau}\}_{i=1}^{n}$. The role of input-output links is limited to amplification; the response of output growth in sector $i$ to a TFP shock in sector $j$ is scaled by element $(i,j)$ of the Leontief-inverse $(I-A)^{-1}$. By summing up connections of all order from $i$ to $j$, the latter shows the importance of sector $j$ in sector $i$’s supply chain.
@@
and for $q\in\mathbb{N}$. Other things equal, under a higher $q$ the economy approaches faster the new steady state level of output following a TFP shock. Relatedly, under lagged IOC, it also means that the IRFs at any single horizon increasingly depend on higher-order connections, and it leads to network-induced cross-sectional correlation in innovations at observational frequency, whereas for $q=1$ the correlation is entirely due to aggregate TFP shocks.474747 See Section C.1 as well as Section 2.2.3 for more detailed discussions of these points. In the limit as $q\rightarrow\infty$, the long-term response referenced in Proposition 6 – a function of all connection-orders – materializes after a single observational period. As the meaning of $q$ depends on the frequency of observation, its choice is discussed in Section 4.2.

### v7

@@ 2.3 Further Properties of the NVAR
Trivially, contemporaneous network interactions lead to contemporaneous correlation among $\{x_{i}\}_{i=1:n}$; under Eq. 6, we have $\mathbb{V}[x]=\mathbb{V}[(I-A)^{-1}v]$. In this case, $\text{Cov}(x_{i},x_{j})$ reflects the bilateral exposure between $i$ and $j$ and their mutual exposure to third units, whereby exposure is determined by network connections of all order. [^21] We have $\mathbb{V}[x]=\mathbb{V}[(I+A+A^{2}+...)v]$, provided that $|\lambda_{i}|<1$ for all eigenvalues $\lambda_{i}$ of $A$. Assuming that $\mathbb{V}[v]=\Sigma=diag(\sigma^{2}_{1},...,\sigma^{2}_{n})$, we get $\text{Cov}(x_{i},x_{j})=\sigma^{2}_{i}\mathbf{1}\left\{{i=j}\right\}+\sigma^{2}_{j}\sum_{h=1}^{\infty}\left[{A^{h}}\right]_{ij}+\sigma^{2}_{i}\sum_{h=1}^{\infty}\left[{A^{h}}\right]_{ji}+\sum_{o=1}^{n}\sigma^{2}_{o}\sum_{h_{i}=1}^{\infty}\sum_{h_{j}=1}^{\infty}\left[{A^{h_{i}}}\right]_{io}\left[{A^{h_{j}}}\right]_{jo}\;.$ De Graeve and Schneider (2023) exploit this insight for shock identification in the context of a production economy.
@@
Corollary 3 above establishes that an SAR for $x$ is obtained in the limit when writing $x$ as the sum of an underlying process driven by lagged network interactions that evolves at an ever higher frequency. In addition, by Proposition 3, the Normality-assumption $v\sim N(0,\Sigma)$ in the SAR model can be justified by temporally independent high-frequency innovations $\tilde{u}_{\tau}$ with $\mathbb{V}[\tilde{u}_{\tau}]=q^{-1}\Sigma$. Concretely, for the NVAR($p,q$) from Eq. 5 and for $a=\sum_{l=1}^{p}\alpha_{l}$, Proposition 3 shows that $\sqrt{q}u_{t}\overset{d}{\rightarrow}N\left({0,\mathbb{V}[(I-aA)^{-1}\tilde{u}_{\tau}]}\right)$ as $q\rightarrow\infty$, provided the high-frequency innovations $\tilde{u}_{\tau}$ are strictly stationary.
@@
The NVAR is a restricted VAR in which innovations transmit across series along constant, bilateral links (in one direction). Relative to a VAR($p$) with $p>1$, the NVAR($p,1$) induces parsimony; e.g. for each $y_{it}$, the $n$ links $\{a_{ij}\}_{j=1:n}$ in $A_{i\cdot}$ summarize the information in the $np$-dimensional set $\{y_{j,t-l}\}_{j=1:n,l=1:p}$:
@@
where $x_{it}=\sum_{j=1}^{n}a_{ij}y_{j,t-l}$ is the $i$th row of $X_{t}=[Ay_{t-1},...,Ay_{t-l}]$. If, in addition, the network $A$ is sparse – as is the case across a wide range of applications1111todo: 11cite some general ref.; Jackson-book? who else? –, then the NVAR rationalizes the dynamic comovement among all $\{y_{it}\}_{i=1:n}$ by the dynamic innovation transmission along few bilateral links among $i=1:n$.
The Dynamic Factor Model (DFM) also reduces dimensionality by summarizing a large set of predictors by a few linear combinations.1212todo: 12cite here; Ambrogio’s paper, and look up original source? Proposition 4 shows that the NVAR($p,1$) can be written as a DFM, while a DFM with restricted factor dynamics can be written as an NVAR($p,1$) in the limit as $n\rightarrow\infty$. [^22] The second part of Proposition 4 restricts $f_{t}\in\mathbb{R}^{r}$ to follow an NVAR($p,1$). This is less restrictive than it may seem; for $p=1$, it requires $f_{t}\sim VAR(p)$, while under $r=1$ it requires $f_{t}\sim AR(p)$. Also, note that, as usual, the factor representation is not unique, and we can re-scale $A$ and $\phi$ to ensure $a_{ij}\in[-1,1]$. Together with Eq. 9 above, it suggests that, at the cost of restricting factor dynamics, the NVAR allows the linear combinations of predictors to vary more flexibly across units, to the point that it naturally acommodates sparse factors as (linear combinations of) locally important units in the network. [^23] The DFM constructs $\{\mathbb{E}_{t-1}[y_{it}]\}_{i=1}^{n}$ as different ($\Lambda_{i\cdot}$) linear combinations of the same factors $f_{t}$, which in turn are linear combinations of observables. In contrast, the NVAR constructs $\{\mathbb{E}_{t-1}[y_{it}]\}_{i=1}^{n}$ as the same ($\alpha$) linear combination of different covariates $x_{it}$, which in turn are linear combinations of observables.
@@ 3.1 Timing of Network Effects
Section B.1 establishes consistency and asymptotic Normality of $\hat{\alpha}_{OLS}$ under large $T$, large $n$ and large $(n,T)$ asymptotics. The derivations under large $n$ assume that the observed network adjacency matrix $A_{n}$ converges to some limit $A_{*}$ so that, for example, $\frac{1}{n}\sum_{i=1}^{n}\left({A_{n,i\cdot}y_{t-l}}\right)^{\prime}u_{it}\overset{p}{\rightarrow}\mathbb{E}\left[{\left({A_{*,i\cdot}y_{t-l}}\right)^{\prime}u_{it}}\right]$. 1313todo: 13verify; at least notation could be better: $y_{t}$ also depends on $n$; it’s not $A$ taht converges, but network-statistics
@@
In case of an NVAR($p,q$), as defined in Eq. 4, an estimator for $\alpha$ can be obtained by data augmentation. [^24] The discussion holds likewise for the NVAR($p,q$) from Eq. 5. However, point identification is not guaranteed. For example, under $q=2$ and $p=1$, the observed process follows
@@
Conditioning on $\tilde{\Sigma}$, the posterior $p(\alpha|Y_{1:T})$ can be obtained using the Gibbs sampler of Carter and Kohn (1994). Treating the unobserved data in $\tilde{Y}_{1:T_{\tau}}$ as parameters, it iteratively draws from $p(\tilde{Y}_{1:T_{\tau}}|\alpha,Y_{1:T})$ and $p(\alpha|\tilde{Y}_{1:T_{\tau}})$ to obtain a sample from $p(\alpha,\tilde{Y}_{1:T_{\tau}}|Y_{1:T})$. [^26] The particular state space model where in some periods $\tau$ no data is observed implies that in these periods the updating-step of the Kalman filter is skipped and the updated distribution of states equals their predicted distribution. See Section B.1, and see Schorfheide and Song (2015) for a discussion of the analogous case of a mixed-frequency VAR. Under a Uniform prior for $\alpha$, the resulting posterior mode of $\alpha$ converges to the ML estimator obtained using the Expectation-Maximization (EM) algorithm.1515todo: 15citation? or show in appendix?! To estimate $\tilde{\Sigma}$ as well, an additional iteration step is added to the Gibbs sampler. Using a uniform prior, we get an Inverse-Wishart conditional posterior for $\Sigma|\tilde{Y}_{1:T_{\tau}},\alpha$ with mode $\frac{1}{T_{\tau}}\sum_{\tau=1}^{T_{\tau}}\tilde{u}_{\tau}\tilde{u}_{\tau}^{\prime}$.
@@ 3.2 Joint Inference: Network & Effect-Timing
To render $(\alpha,A)$ jointly identified, I normalize $\alpha_{1}=1$, with appropriate redefinitions of $\alpha$ as well as $y_{t}$, $z_{t}$ and $X_{t}$. Relative to the alternative of restricting the norm of $\alpha$, this normalization facilitates analytical calculations and asymptotic analysis, but it requires $\alpha_{1}\neq 0$ in the true data generating process. [^27] In case $a_{ij}\geq 0$ is restricted, it requires $\alpha_{1}>0$.
@@
Long and Plosser (1983) show that time lags between the production of goods and their subsequent use as intermediaries for producing other goods can generate endogenous business cycles. In this application, I empirically quantify the importance of their proposed channel. I generalize their RBC economy with one-period lagged IOC by assuming that firms’ production requires inputs produced in the past $p$ periods. This leads to sectoral output growth evolving at some model-frequency as an NVAR($p,1$), which translates into an NVAR($p,q$) for some $q\in\mathbb{N}$ at my monthly frequency of observation. Thereby, $A$ is the input-output matrix, $u_{t}$ contains sectoral productivity processes, while $\{\alpha_{l}\}_{l=1:p}$ show how input-sourcing is spread out over the $p$ periods. By estimating $\alpha$, $q$ and the persistence in $u_{t}$, I use the NVAR to quantify the extent to which business cycles in this framework are due to lagged IOC as opposed to persistence in exogenous productivity processes.
After theoretically motivating the analysis in Section 4.1, I discuss the setup, data and estimation in Section 4.2. Section 4.3 presents the results. Details are in Appendix C.
@@ 4.1 Theory
The following analysis is based on Carvalho and Tahbaz-Salehi (2019) – who discuss a static economy – and the Appendix to Acemoglu et al. (2016). Derivations are in Section C.1.
@@
where $\sum_{i=1}^{n}\gamma_{i}=1$. The first-order condition (FOC) yields $c_{i\tau}=\gamma_{i}\frac{w_{\tau}}{p_{i\tau}}$. [^30] Hence, $\gamma_{i}$ is the share of good $i$ in households’ expenditures. This result holds even if households have access to a storage technology, as market clearing under representative households in a closed economy implies that households spend their whole period $\tau$ income, $w_{\tau}$, on consumption. 1717todo: 17see P009diary3, p.152 for dynamic HH problem
@@
where $k^{y}$ is a vector of constants, and $a_{ij}=p_{j}x^{ij}/(p_{i}y_{i})$ is value of good $j$ purchases by sector $i$ as a fraction of the value of sector $i$’s output. In this environment, while the network $A$ amplifies idiosyncratic TFP shocks and therefore affects the variance of $\tilde{y}$, any autocorrelation in $\tilde{y}$ is inherited from that in $\tilde{z}$ (see e.g. Carvalho (2010); Acemoglu et al. (2012)).
To analyze dynamics under lagged IOC, I assume perfect foresight. [^31] As discussed in Fan et al. (2023), this assumption is standard for modeling dynamic spatial economies, which are closely related to dynamic network economies. If, as in Long and Plosser (1983), it takes one period to convert purchased inputs into output, then $x_{ij\tau}=x^{ij}_{\tau-1}$ and sectoral output approximately follows an NVAR($1,1$):
@@
where the time variation in $k^{y1}_{\tau}$ is due to time variation in the numéraire $w_{\tau}$. So long as $\beta<1$, in steady state, prices are higher and output is lower than in the economy with contemporaneous IOC. Also, the meaning of $a_{ij}=\beta^{-1}p_{j}x^{ij}/(p_{i}y_{i})$ changes slightly. These differences vanish as $\beta\rightarrow 1$. More importantly, while the former, static economy is always in steady state, the economy with one period-lagged IOC features transition dynamics; after a disturbance to $\tilde{z}_{\tau}$, $\tilde{y}_{\tau}$ only asymptotically converges to steady state. [^32] Relatedly, as discussed in Section 2.2, the response of $\tilde{y}_{(\tau)}$ to a change in $\tilde{z}_{(\tau)}$ in the static economy corresponds to the long-term response of $\tilde{y}_{\tau}$ to a permanent change in $\tilde{z}_{\tau}$ in this dynamic economy (disregarding the slightly changed meaning of $a_{ij}$). It is due to these transition dynamics that this framework can generate endogenous business cycles, i.e. autocorrelation in $\tilde{y}_{\tau}$ even in absence of autocorrelation in $\tilde{z}_{\tau}$.
@@
where $x^{ij}_{\tau,\tau-h}$ denotes the use of good $j$ purchased at time $\tau-h$ in the production of good $i$ at time $\tau$. [^33] This means that a good perishes after two periods (with regard to its suitability as an input in production). Therefore, the amount of good $j$ purchased at time $\tau$ can be used in production at periods $\tau+1$ and $\tau+2$: $x^{ij}_{\tau}=x^{ij}_{\tau+1,\tau}+x^{ij}_{\tau+2,\tau}$. This shortcut stands for frictions like delivery costs and -lags and storage capacity constraints (Khan and Thomas, 2007; Alessandria et al., 2010; Liu and Tsyvinski, 2024; Antràs and Tubdenov, 2025). [^34] As in the Long and Plosser (1983)-economy above, the presumption is that storage is done by the buyer. In the Cobb-Douglas case $r\rightarrow 0$, sectoral output approximately follows an NVAR(2,1):
@@
where once again $k^{y2}_{\tau}$ varies over time only to the extent that the numéraire $w_{\tau}$ changes in value. Under a more general elasticity of substitution $r\in[0,1)$, [^35] This notably excludes complementarity ($r<0$) and perfect substitutability ($r=1$). the analogous result is obtained by log-linearizing around the steady state (see Section C.1). This specification nests the one period-lagged economy, which is obtained under $\alpha_{2}=0$. Relative to that case, $\alpha_{2}>0$ increases prices and decreases output in steady state, provided that $\beta<1$. Also, we have $a_{ij}=\left[{\beta\alpha_{1}+\beta^{2}\alpha_{2}}\right]^{-1}(p_{j}x^{ij})/(p_{i}y_{i})$. As $\beta\rightarrow 1$, the links $a_{ij}$ retain their interpretation as the output shares of different inputs $j$ in the production of good $i$. The parameters $(\alpha_{1},\alpha_{2})$ show the shares of an input $j$ purchased at different periods in the past in the overall usage of input $j$ in the production of good $i$. Their homogeneity means that the time profile of input-sourcing is assumed to be constant over time and across input-output pairs $(i,j)$. The restrictions $\alpha_{1},\alpha_{2}\geq 0$, $\alpha_{1}+\alpha_{2}=1$ and $\sum_{j=1}^{n}a_{ij}<1\;\forall\;i$ imply that $\tilde{y}_{\tau}$ is stationary as long as $\tilde{z}_{\tau}$ is. [^36] Berman and Plemmons (1994, ch. 2) show that for an element-wise nonnegative matrix with row sums strictly smaller than 1, the absolute value of the largest eigenvalue is strictly less than 1. Stationarity then follows by Corollary 1.
The theory above does not restrict the process of (log) sectoral TFP $\tilde{z}_{i\tau}$. In general, it may display secular growth and persistent shocks of both aggregate and idioyncratic nature. Consider a difference-stationary specification with aggregate and idiosyncratic TFP processes $e^{a}_{\tau}$ and $e_{i\tau}$ evolving as AR(1) processes as in Foerster et al. (2011):
@@
We obtain IRFs $\partial\Delta\tilde{y}_{i,\tau+h}/\partial\varepsilon_{j,\tau}$ and $\partial\Delta\tilde{y}_{i,\tau+h}/\partial\varepsilon^{a}_{\tau}$ akin to those in Eq. 7 in Section 2.3. This points to two sources of persistence in output growth: persistence in the aggregate TFP process $e^{a}_{\tau}$ and persistence in idiosyncratic TFP processes $\{e_{i\tau}\}_{i=1}^{n}$. The role of input-output links is limited to amplification; the response of output growth in sector $i$ to a TFP shock in sector $j$ is scaled by element $(i,j)$ of the Leontief-inverse $(I-A)^{-1}$, as in Carvalho (2010, Fig. 7 and 8). By summing up connections of all order from $i$ to $j$, the latter shows the importance of sector $j$ in sector $i$’s supply chain.
@@
and for $q\in\mathbb{N}$. Other things equal, under a higher $q$ the economy approaches faster the new steady state level of output following a TFP shock. Relatedly, under lagged IOC, it also means that the IRFs at any single horizon increasingly depend on higher-order connections, and it leads to network-induced cross-sectional correlation in innovations at observational frequency, whereas for $q=1$ the correlation is entirely due to aggregate TFP shocks. [^37] See Section C.1 as well as Section 2.2.3 for more detailed discussions of these points. In the limit as $q\rightarrow\infty$, the long-term response referenced in Proposition 6 – a function of all connection-orders – materializes after a single observational period. As the meaning of $q$ depends on the frequency of observation, its choice is discussed in Section 4.2.